OpenStudy (anonymous):

[Reward : Medal, Fan and Friendship :3] Factorise m^4 - 16 completely using the suitable identity.

OpenStudy (anonymous):

@ganeshie8 @Iambatman

OpenStudy (anonymous):

For this you need to have: \[\large \color{red}{a^2 - b^2 = (a+b)(a-b)}\]

OpenStudy (anonymous):

See, you can rearrange what you are given with like the following: \[\color{blue}{m^4 = (m^2)^2}\] \[\color{green}{16 = 4^2}\]

OpenStudy (anonymous):

So, now make use of that formula: \[(\color{blue}{m^2})^2 - (\color{green}{4})^2 = ??\]

OpenStudy (anonymous):

So, basically, you are having : \(a = m^2\) and \(b = 4\). Just plug in these values of \(a\) and \(b\) there in that formula and there you go..!!

OpenStudy (anonymous):

Do you want to try yourself or just want answer?

OpenStudy (anonymous):

Can I have the answer coz Im kinda busy? @eta

OpenStudy (anonymous):

Yes, we are aliens here who are totally free to do your homework, right?? Do the problem yourself, I have given sufficient data to solve it, and please try once and if having doubt, then do ask me..

OpenStudy (anonymous):

Thanks I guess. -_- @eta

OpenStudy (anonymous):

Don't guess, make yourself sure everytime.. :P

OpenStudy (anonymous):

Cheese, Fine then Thanks. :P

OpenStudy (anonymous):

\(\dagger\)